# The Relative Entropy of Expectation and Price

### v6

@@ III.3 Market and model risk
00 0 $=\mathbb{E}^{\phi}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(P-\delta\cdot dq)]]$
@@
are solved to generate the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$ for the derivative security. While the hedge neutralises the underlying component of the derivative risk, when the market is incomplete there are unhedged risks that are sensitive to the choice of expectation measure. The entropic pricing framework distinguishes between these market and model risks in the derivative security.
Consider an alternative choice $\bar{\mathbb{E}}$ for the expectation measure, related to the original expectation measure $\mathbb{E}$ by the Radon-Nikodym weight:
$\frac{d\bar{\mathbb{E}}}{d\mathbb{E}}=\frac{\displaystyle\exp[\epsilon W]}{\displaystyle\mathbb{E}[\exp[\epsilon W]]}=1+\epsilon(W-\mathbb{E}[W])+O[\epsilon^{2}]$ (66)
for a bounded weight $W$ and a scale parameter $\epsilon$. Using this alternative expectation measure in the calibration conditions, the implied funding rate and unit optimal portfolio are adjusted, $\bar{r}=r+\epsilon\,\partial r$ and $\bar{\phi}=\phi+\epsilon\,\partial\phi$, with:
$\partial r$ $=\frac{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,W]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}+O[\epsilon]$ (67)
@@
Applying these adjustments, the original and alternative price measures are then related by:
@@
Similar adjustments apply to the implied funding spread, hedge portfolio and initial price, derived from the hedge and price conditions in the alternative expectation measure:
@@
00 0 $=\bar{\mathbb{E}}^{\bar{\phi}}[(dq-q(\bar{r}+\alpha\bar{s})\,dt)\exp[-\alpha(P-\bar{\delta}\cdot dq)]]$
@@
IV Markovian diffusion
In this section, assume that there is a vector $x$ of state variables whose cumulant generating function in the expectation measure $\mathbb{E}$ satisfies:
@@ IV Markovian diffusion
As a concrete example of a continuous Markovian diffusion, the Heston model illustrates the key features of the more general framework. In this model, the state comprises a funding price $b$, an underlying price $q$ and a variance $\nu$ satisfying the stochastic differential equations:
@@
$d\nu$ $=(\bar{\nu}-\nu)\kappa\,dt+\sqrt{\nu}\,\sigma\,dz$
where $dy$ and $dz$ are Brownian increments in the expectation measure with correlation $\rho$. This is equivalent to the cumulant generating function:
@@
$\beta\,br+\psi\,q\mu+\eta\,(\bar{\nu}-\nu)\kappa$
$+\frac{1}{2}\psi^{2}\,q^{2}\nu+\psi\eta\,q\nu\rho\sigma+\frac{1}{2}\eta^{2}\,\nu\sigma^{2}+O[dt]$
The funding price has predictable increment with mean $br\,dt$ where $r$ is the funding rate. The underlying price has increment with mean $q\mu\,dt$ and variance $q^{2}\nu\,dt$ where $\mu$ is the mean rate. The variance has increment with mean $(\bar{\nu}-\nu)\kappa\,dt$ and variance $\nu\sigma^{2}\,dt$ where $\kappa$ is the mean reversion rate, $\bar{\nu}$ is the mean reversion level and $\sigma$ is the volatility of variance. The final parameter $\rho$ is the correlation between the increments of the underlying price and the variance.
Since the funding security has predictable return, the implied funding rate is $r$ and the implied funding spread is $s=0$. In order to satisfy the self-funding condition, the portfolio with underlying weight $\phi$ must have funding weight $-\phi q/b$, and the cumulant generating function in the price measure $\mathbb{E}^{\phi}$ satisfies:
@@
$\beta\,br+\psi\,q\mu^{\phi}+\eta\,(\bar{\nu}^{\phi}-\nu)\kappa$
$+\frac{1}{2}\psi^{2}\,q^{2}\nu+\psi\eta\,q\nu\rho\sigma+\frac{1}{2}\eta^{2}\,\nu\sigma^{2}+O[dt]$
where the mean rate of the underlying price and the mean reversion level of the variance are adjusted:
@@
$\bar{\nu}^{\phi}$ $:=\bar{\nu}-\phi q\nu\frac{\rho\sigma}{\kappa}$
@@
$\bar{\nu}^{\phi}$ $=\bar{\nu}-(\mu-r)\frac{\rho\sigma}{\kappa}$
@@
$\delta={}$ $\frac{\partial p}{\partial q}+\frac{\partial p}{\partial\nu}\frac{\rho\sigma}{q}+O[dt]$ (86)
$0={}$ $\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial b}br+\frac{\partial p}{\partial q}qr+\frac{\partial p}{\partial\nu}((\bar{\nu}-\nu)\kappa-(\mu-r)\rho\sigma)$
$+\frac{1}{2}\frac{\partial^{2}p}{\partial q^{2}}\,q^{2}\nu+\frac{\partial^{2}p}{\partial q\,\partial\nu}\,q\nu\rho\sigma+\frac{1}{2}\frac{\partial^{2}p}{\partial\nu^{2}}\,\nu\sigma^{2}$
$-\frac{1}{2}\alpha\left(\frac{\partial p}{\partial\nu}\right)^{\!\!2}\!\nu\sigma^{2}(1-\rho^{2})+O[dt]$
for the hedge portfolio $\delta[t,b,q,\nu]$ and the derivative price $p[t,b,q,\nu]$. In the first expression, the hedge ratio is adjusted when the increment of the variance is correlated with the increment of the underlying price. The derivative price then satisfies a linear parabolic differential equation with a non-linear correction term that prices the unhedgeable risks when the increments are not perfectly correlated.
V Discrete models
@@
Distribution ($n=$100$$) Distribution ($n=$1000$$) Distribution ($n=$10000$$)
Figure 8: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=$100$$, $1000$ and $10000$ samples from a standardised normal variable, centralised to have zero mean. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price $q=-0.4$ (top row), $q=0$ (middle row) and $q=0.4$ (bottom row). In the transformation from expectation to price measure, the uniform probabilities are biased using an exponential function of the final underlying price. Among equivalent measures that satisfy the calibration condition, the price measure thus has minimum entropy relative to the expectation measure.
@@
Call Option ($n=$100$$) Call Option ($n=$1000$$) Call Option ($n=$10000$$)
@@
Digital Option ($n=$100$$) Digital Option ($n=$1000$$) Digital Option ($n=$10000$$)
Figure 9: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=$100$$, $1000$ and $10000$ samples from a standardised normal variable. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price $q=-0.4$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Call Option ($n=$100$$) Call Option ($n=$1000$$) Call Option ($n=$10000$$)
@@
Digital Option ($n=$100$$) Digital Option ($n=$1000$$) Digital Option ($n=$10000$$)
Figure 10: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=$100$$, $1000$ and $10000$ samples from a standardised normal variable. The price measure is supported on the same samples with uniform probabilities that match the target initial underlying price $q=0$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Call Option ($n=$100$$) Call Option ($n=$1000$$) Call Option ($n=$10000$$)
@@
Digital Option ($n=$100$$) Digital Option ($n=$1000$$) Digital Option ($n=$10000$$)
Figure 11: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on $n=$100$$, $1000$ and $10000$ samples from a standardised normal variable. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price $q=0.4$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Figure 12: In these studies of discrete models on quantum information, the final hedging price is uniformly distributed in the expectation measure on $n=2$ values $-1$ and $1$, and is decorrelated from the final funding price by applying the rotation $\theta=0$, $\pi/5$ and $\pi/4$ to its eigenbasis. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
Entropy is a universal concept in information theory that is ideally suited to a wide range of applications. In this section, the entropic risk metric is applied to risk optimisation in discrete models of classical and quantum information. The core premise of the discrete model is that the indeterminate variable has a spectrum of possible eigenvalues, which may include duplicates, that are uniformly distributed in the expectation measure. Classical or quantum models are then characterised by the existence or absence of a common basis of eigenstates supporting the eigenvalues.
@@
00 0 $=\operatorname{tr}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (90)
00 0 $=\operatorname{tr}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]$
@@
These conditions are explored in the studies of figures 9, 10, 11 and 12, assuming an underlying market that contains a funding security with initial price $b=1$ and final price $B=b+db$ satisfying $\operatorname{tr}[db]=0$ and a hedging security with initial price $q$ and final price $Q=q+dq$ satisfying $\operatorname{tr}[Q]=0$. The model is developed for the final derivative price expressed as a function $P=f[Q-kB]$ of the spread $Q-kB$ at strike $k$.
V.1 Classical study
The classical studies of figures 9, 10 and 11 assume $B=1$ and $Q$ is a diagonal matrix with eigenvalues $Q_{i}$ sampled from a standard normal variable, centralised to have zero sample mean. The matrix $P=f[Q-kB]$ is then diagonal with eigenvalues $P_{i}:=f[Q_{i}-k]$.

### v7

@@ IV Market and model risk
$0$ $=\mathbb{E}^{\phi}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(P-\delta\cdot dq)]]$
@@
are solved to generate the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$ for the derivative security. While the hedge neutralises the underlying component of the derivative risk, when the market is incomplete there are unhedged risks that are sensitive to the choice of expectation measure. This is the origin of model risk.
Suppose that the after expectation measure $\bar{\mathbb{E}}$ is equivalent to the before expectation measure $\mathbb{E}$ with Radon-Nikodym weight:
$\frac{\displaystyle d\bar{\mathbb{E}}}{\displaystyle d\mathbb{E}}=\frac{\displaystyle\exp[\epsilon W]}{\displaystyle\mathbb{E}[\exp[\epsilon W]]}=1+\epsilon(W-\mathbb{E}[W])+O[\epsilon^{2}]$ (112)
for a bounded weight $W$ and a scale parameter $\epsilon$. With the benefit of hindsight, the implied funding rate and unit optimal portfolio are adjusted, $\bar{r}=r+\epsilon\,\partial r$ and $\bar{\phi}=\phi+\epsilon\,\partial\phi$, with:
$\partial r$ $=\frac{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,W]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\cdot q\,dt}+O[\epsilon]$ (113)
@@
Applying these adjustments, the before and after price measures are related by:
@@
Similar adjustments apply to the implied funding spread, hedge portfolio and initial price, derived from the hedge and price conditions in the after expectation measure:
@@
$0$ $=\bar{\mathbb{E}}^{\bar{\phi}}[(dq-q(\bar{r}+\alpha\bar{s})\,dt)\exp[-\alpha(P-\bar{\delta}\cdot dq)]]$
@@
V The Black-Scholes model
As an illustration of the entropic method, consider an underlying market comprising a funding security with price $b>0$ and an underlying security with price $q$ satisfying the stochastic differential equations:
$db$ $=br\,dt$ (117)
$dq$ $=q(\mu\,dt+\sqrt{\nu}\,dy)$
where the increment $dy$ is Brownian in the expectation measure. The model is parametrised by the funding rate $r$ of the funding security and the mean rate $\mu$ and variance rate $\nu>0$ of the underlying security.
The self-funded portfolio in the underlying market comprises $\omega$ units of the underlying security funded by $\omega(q/b)$ units of the funding security. The portfolio return has entropy-adjusted mean:
$\frac{1}{dt}\mathbb{E}[$ $\alpha][\omega(dq-(q/b)db)]=$ (118)
$\frac{1}{2}\frac{(\mu-r)^{2}}{\alpha\nu}-\frac{1}{2}\alpha q^{2}\nu\left(\omega-\frac{\mu-r}{\alpha q\nu}\right)^{\!2}+O[dt]$
This is maximised by the optimal portfolio:
$\omega=\frac{\mu-r}{\alpha q\nu}$ (119)
When the derivative security is added, the self-funded portfolio comprises one unit of the derivative security and $(\omega-\delta)$ units of the underlying security funded by $(p/b)-(\omega-\delta)(q/b)$ units of the funding security. The underlying market is complete when the model parameters are fixed but incomplete when the model parameters are volatile. In the first scenario, risk can be neutralised by hedging and the derivative price is determined from the no-arbitrage condition. In the second scenario, risk cannot be neutralised by hedging and the no-arbitrage condition does not determine the derivative price. By replacing risk neutralisation with risk optimisation as the objective of hedging, the entropic method extends to this case. The derivative price is then resolved by replacing no-arbitrage with indifference as the equilibrium condition for pricing.
Complete market
Assume the funding rate $r$, mean rate $\mu$ and variance rate $\nu$ are fixed. The portfolio return has entropy-adjusted mean:
$\frac{1}{dt}\mathbb{E}[$ $\alpha][(dp-(p/b)db)+(\omega-\delta)(dq-(q/b)db)]=$ (120)
$\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial q}qr+\frac{1}{2}\frac{\partial^{2}p}{\partial q^{2}}q^{2}\nu+\frac{1}{2}\frac{(\mu-r)^{2}}{\alpha\nu}$
$-\frac{1}{2}\alpha q^{2}\nu\left((\omega-\delta)-\left(\frac{\mu-r}{\alpha q\nu}-\frac{\partial p}{\partial q}\right)\right)^{\!2}+O[dt]$
when the derivative security with price $p[t,q]$ is added. This is maximised by the hedge portfolio:
$\delta=\frac{\partial p}{\partial q}$ (121)
The indifference condition dictates that the addition of the derivative security neither increases nor decreases the maximum entropy-adjusted mean of the portfolio, leading to the equation:
$0=\frac{\partial p}{\partial t}-(p-\delta q)r+\frac{1}{2}\frac{\partial^{2}p}{\partial q^{2}}q^{2}\nu$ (122)
for the derivative price. The expression on the right is the sum of time decay, funding and gamma contributions to the profit and loss of the hedged derivative. Setting this to zero eliminates the possibility of arbitrage.
Incomplete market
Assume the funding rate $r$ and mean rate $\mu$ are fixed and the variance rate $\nu$ satisfies the stochastic differential equation:
$d\nu=\gamma[t,q,\nu]\,dt+\epsilon[t,q,\nu]\sqrt{\nu}\,dz$ (123)
where the increments $dy$ and $dz$ are joint Brownian with correlation $\rho[t,q,\nu]$ in the expectation measure. The portfolio return has entropy-adjusted mean:
$\frac{1}{dt}\mathbb{E}[$ $\alpha][(dp-(p/b)db)+(\omega-\delta)(dq-(q/b)db)]=$ (124)
$\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial q}qr+\frac{\partial p}{\partial\nu}(\gamma-(\mu-r)\rho\epsilon)$
$+\frac{1}{2}\frac{\partial^{2}p}{\partial q^{2}}q^{2}\nu+\frac{\partial^{2}p}{\partial q\partial\nu}q\nu\rho\epsilon+\frac{1}{2}\frac{\partial^{2}p}{\partial\nu^{2}}\nu\epsilon^{2}$
$-\frac{1}{2}\alpha\left(\frac{\partial p}{\partial\nu}\right)^{\!2}\!\nu(1-\rho^{2})\epsilon^{2}+\frac{1}{2}\frac{(\mu-r)^{2}}{\alpha\nu}$
$-\frac{1}{2}\alpha q^{2}\nu\left((\omega-\delta)-\left(\frac{\mu-r}{\alpha q\nu}-\frac{\partial p}{\partial q}-\frac{\partial p}{\partial\nu}\frac{\rho\epsilon}{q}\right)\right)^{\!2}$
$+O[dt]$
when the derivative security with price $p[t,q,\nu]$ is added. This is maximised by the hedge portfolio:
$\delta=\frac{\partial p}{\partial q}+\frac{\partial p}{\partial\nu}\frac{\rho\epsilon}{q}$ (125)
The indifference condition dictates that the addition of the derivative security neither increases nor decreases the maximum entropy-adjusted mean of the portfolio, leading to the equation:
$0={}$ $\frac{\partial p}{\partial t}-(p-\delta q)r+\frac{\partial p}{\partial\nu}(\gamma-\mu\rho\epsilon)$ (126)
$+\frac{1}{2}\frac{\partial^{2}p}{\partial q^{2}}q^{2}\nu+\frac{\partial^{2}p}{\partial q\partial\nu}q\nu\rho\epsilon+\frac{1}{2}\frac{\partial^{2}p}{\partial\nu^{2}}\nu\epsilon^{2}$
$-\frac{1}{2}\alpha\left(\frac{\partial p}{\partial\nu}\right)^{\!2}\!\nu(1-\rho^{2})\epsilon^{2}$
for the derivative price. The expression on the right is the sum of time decay, funding, vega, gamma and value-at-risk contributions in a conservative estimate of the profit and loss of the hedged derivative. Setting this to zero generates the fair price of the derivative security, including a reserve for the residual risk that remains after hedging.
VI Markovian diffusion
Generalising the example of the Black-Scholes model, in this section assume that there is a vector $x$ of state variables whose cumulant generating function in the expectation measure $\mathbb{E}$ satisfies:
@@ VI Markovian diffusion
As a concrete example of a continuous Markovian diffusion, the Heston model illustrates the key features of the more general framework. In this model, the state variables are the funding price $b$, the underlying price $q$ and the variance rate $\nu$ satisfying the stochastic differential equations:
@@
$d\nu$ $=\kappa(\bar{\nu}-\nu)\,dt+\epsilon\sqrt{\nu}\,dz$
where $dy$ and $dz$ are joint-Brownian increments in the expectation measure with correlation $\rho$. This is equivalent to the cumulant generating function:
@@
$\beta\,br+\psi\,q\mu+\eta\,\kappa(\bar{\nu}-\nu)$
$+\frac{1}{2}\psi^{2}\,q^{2}\nu+\psi\eta\,q\nu\rho\epsilon+\frac{1}{2}\eta^{2}\,\nu\epsilon^{2}+O[dt]$
The funding price has predictable increment with mean brdtbr\,dt where $r$ is the funding rate. The underlying price has increment with mean $q\mu\,dt$ and variance $q^{2}\nu\,dt$ where $\mu$ is the mean rate. The variance rate has increment with mean $\kappa(\bar{\nu}-\nu)\,dt$ and variance $\nu\epsilon^{2}\,dt$ where $\kappa$ is the mean reversion rate, $\bar{\nu}$ is the mean reversion level and $\epsilon$ is the volatility of variance. The final parameter $\rho$ is the correlation between the increments of the underlying price and the variance rate.
Since the funding security has predictable return, the implied funding rate is $r$ and the implied funding spread is $s=0$. The self-funded portfolio with underlying weight $\phi$ and funding weight $\phi(q/b)$ creates the price measure $\mathbb{E}^{\phi}$ with cumulant generating function:
@@
$\beta\,br+\psi\,q\mu^{\phi}+\eta\,\kappa(\bar{\nu}^{\phi}-\nu)$
$+\frac{1}{2}\psi^{2}\,q^{2}\nu+\psi\eta\,q\nu\rho\epsilon+\frac{1}{2}\eta^{2}\,\nu\epsilon^{2}+O[dt]$
where the mean rate of the underlying price and the mean reversion level of the variance rate are adjusted:
@@
$\bar{\nu}^{\phi}$ $:=\bar{\nu}-\phi q\nu\frac{\rho\epsilon}{\kappa}$
@@
$\bar{\nu}^{\phi}$ $=\bar{\nu}-(\mu-r)\frac{\rho\epsilon}{\kappa}$
@@
$\delta={}$ $\frac{\partial p}{\partial q}+\frac{\partial p}{\partial\nu}\frac{\rho\epsilon}{q}+O[dt]$ (142)
$0={}$ $\frac{\partial p}{\partial t}-pr+\frac{\partial p}{\partial q}qr+\frac{\partial p}{\partial\nu}(\kappa(\bar{\nu}-\nu)-(\mu-r)\rho\epsilon)$
$+\frac{1}{2}\frac{\partial^{2}p}{\partial q^{2}}\,q^{2}\nu+\frac{\partial^{2}p}{\partial q\,\partial\nu}\,q\nu\rho\epsilon+\frac{1}{2}\frac{\partial^{2}p}{\partial\nu^{2}}\,\nu\epsilon^{2}$
$-\frac{1}{2}\alpha\left(\frac{\partial p}{\partial\nu}\right)^{\!2}\!\nu(1-\rho^{2})\epsilon^{2}+O[dt]$
for the hedge portfolio $\delta[t,q,\nu]$ and the derivative price $p[t,q,\nu]$. In the first expression, the hedge ratio is adjusted when the increment of the variance rate is correlated with the increment of the underlying price. The derivative price then satisfies a linear parabolic differential equation with a non-linear correction term that prices the unhedgeable risks when the increments are not perfectly correlated.
VII Discrete models
@@
Distribution (n=100n=\$100\$) Distribution (n=1000n=\$1000\$) Distribution (n=10000n=\$10000\$)
Figure 10: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on n=100n=\$100\$, 10001000 and 1000010000 samples from a standardised normal variable, centralised to have zero mean. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price q=0.4q=-0.4 (top row), $q=0$ (middle row) and $q=0.4$ (bottom row). In the transformation from expectation to price measure, the uniform probabilities are biased using an exponential function of the final underlying price. Among equivalent measures that satisfy the calibration condition, the price measure thus has minimum entropy relative to the expectation measure.
@@
Call Option (n=100n=\$100\$) Call Option (n=1000n=\$1000\$) Call Option (n=10000n=\$10000\$)
@@
Digital Option (n=100n=\$100\$) Digital Option (n=1000n=\$1000\$) Digital Option (n=10000n=\$10000\$)
Figure 11: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on n=100n=\$100\$, 10001000 and 1000010000 samples from a standardised normal variable. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price q=0.4q=-0.4. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Call Option (n=100n=\$100\$) Call Option (n=1000n=\$1000\$) Call Option (n=10000n=\$10000\$)
@@
Digital Option (n=100n=\$100\$) Digital Option (n=1000n=\$1000\$) Digital Option (n=10000n=\$10000\$)
Figure 12: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on n=100n=\$100\$, 10001000 and 1000010000 samples from a standardised normal variable. The price measure is supported on the same samples with uniform probabilities that match the target initial underlying price $q=0$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Call Option (n=100n=\$100\$) Call Option (n=1000n=\$1000\$) Call Option (n=10000n=\$10000\$)
@@
Digital Option (n=100n=\$100\$) Digital Option (n=1000n=\$1000\$) Digital Option (n=10000n=\$10000\$)
Figure 13: In these studies of discrete models on classical information, the final underlying price is uniformly distributed in the expectation measure on n=100n=\$100\$, 10001000 and 1000010000 samples from a standardised normal variable. The price measure is supported on the same samples with probabilities adjusted to match the target initial underlying price $q=0.4$. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
@@
Figure 14: In these studies of discrete models on quantum information, the final hedging price is uniformly distributed in the expectation measure on $n=2$ values 1-1 and $1$, and is decorrelated from the final funding price by applying the rotation $\theta=0$, $\pi/5$ and $\pi/4$ to its eigenbasis. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
Entropy is a universal concept in information theory that is ideally suited to a wide range of applications. In this section, the entropic risk metric is applied to risk optimisation in discrete models of classical and quantum information. The core premise of the discrete model is that the indeterminate variable has a spectrum of possible eigenvalues, which may include duplicates, that are uniformly distributed in the expectation measure. Classical or quantum models are then characterised by the existence or absence of a common basis of eigenstates supporting the eigenvalues of all the variables.
@@
$0$ $=\operatorname{tr}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (146)
$0$ $=\operatorname{tr}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]$
@@
These conditions are explored in the studies of figures˜11, 12, 13 and 14, assuming an underlying market that contains a funding security with initial price $b=1$ and final price $B=b+db$ satisfying $\operatorname{tr}[db]=0$ and a hedging security with initial price $q$ and final price $Q=q+dq$ satisfying $\operatorname{tr}[Q]=0$. The model is developed for the final derivative price expressed as a function P=f[QkB]P=f[Q-kB] of the spread QkBQ-kB at strike $k$.
VII.1 Classical study
The classical studies of figures˜11, 12 and 13 assume $B=1$ and $Q$ is a diagonal matrix with eigenvalues $Q_{i}$ sampled from a standard normal variable, centralised to have zero sample mean. The matrix P=f[QkB]P=f[Q-kB] is then diagonal with eigenvalues $P_{i}:=f[Q_{i}-k]$.